Mr. Wilson is 15 years older than Mr. Connors. Five years from now, Mr. Wilson will be 1 1/2 times as old as Mr. Connors will be then. How old is each man now?
Found 2 solutions by edjones, TheWastedWizard:
Answer by edjones(8007)

w=c+15
w=1.5c
1.5c=c+15
.5c=15
c=30
w=45
.
c=30-5=25 yo now
w=45-5=40 now.
.
Ed

Answer by TheWastedWizard(1)

Let c be Mr. Connors' age NOW.
In 15 years:
Mr. Wilson will be 20 + c years old
Mr. Connors will be c + 5 years old
Equation:
20 + c = 1.5 (c + 5)
20 + c = 1.5c + 7.5
12.5 + c = 1.5 c
12.5 = 0.5c
25 years = c
Mr. Connors is 25 years old NOW.
Mr. Wilson is 15 years older than Mr. Connors, so he is 15 + 25 years.
Mr. Wilson is 40 years old NOW.